### v2

# Risk-Neutral Pricing Model of Uniswap Liquidity Providing Position: A Stopping Time Approach

Abstract
In this paper, we introduce a novel pricing model for Uniswap V3, built upon stochastic processes and the Martingale Stopping Theorem. This model innovatively frames the valuation of positions within Uniswap V3. We further conduct a numerical analysis and examine the sensitivities through Greek risk measures to elucidate the model’s implications. The results underscore the model’s significant academic contribution and its practical applicability for Uniswap liquidity providers (LPs), particularly in assessing risk exposure and guiding hedging strategies.
Index Terms:
Uniswap, pricing model, greeks, perpetual option, risk management, hedging strategies
I Introduction
Uniswap[1], launched in November 2018, represents a decentralized, open-source protocol for cryptocurrency exchange. As the first automated market maker (AMM) to achieve widespread adoption, Uniswap enables users to conduct token swaps directly from their wallets via liquidity pools. In these pools, liquidity providers (LPs) earn fees by contributing their assets.
The introduction of Uniswap V3 brought significant advancements, most notably the concept of concentrated liquidity. This feature allows LPs to allocate their capital to specific price ranges, enhancing capital efficiency and yielding higher returns while simultaneously mitigating price risk. Furthermore, Uniswap V3 incorporates multiple fee tiers, thereby providing LPs with flexibility in selecting their compensation levels in accordance with their risk tolerance.
With the ongoing expansion of decentralized finance (DeFi) adoption, the demands on LPs in Uniswap v3 have intensified. LPs are now required to meticulously balance a myriad of factors, including risk, returns, and volatility, necessitating a sophisticated understanding and strategic approach to liquidity provision.
II Related Work and Limitations
II-A Loss-Versus-Rebalancing (LVR)
LVR is a metric designed to quantify the impact of adverse selection on automated market maker (AMM) LPs[2]. LVR captures the costs incurred by LPs due to stale prices that better-informed arbitrageurs exploit. Unlike impermanent loss, which only considers the starting and ending points of a trade, LVR compares the performance of LPs to a rebalancing portfolio, highlighting additional costs associated with providing liquidity.
LVR incorporates volatility into its calculations, making it a useful tool for selecting the optimal liquidity pool. However, it remains primarily a risk assessment tool. It does not evaluate the returns for LPs when exiting positions at different prices, nor does it provide guidance on the optimal range for market-making activities.
II-B Guillaume Lambert
Guillaume Lambert focused on pricing within the Black-Scholes-Merton (BSM) framework in his work[3][4]. He viewed liquidity positions as perpetual options and attempted to replicate liquidity positions with covered calls. He also extends his model towards implied volatility and hedging. However, a limitation of Lambert’s model is the specification of an expiration time T, which inherently makes it an approximation rather than a precise solution. Additionally, Lambert explored the use of stochastic processes to calculate the probability of prices remaining within a position range, which has greatly inspired our research. We aim to build upon this foundation by starting from stochastic processes to initiate our study.
II-C Martingale and first hitting problem with two boundaries
Unlike traditional derivatives with specific expiration dates, Uniswap V3 contracts are perpetual, persisting until the price breaches predefined boundaries or the contract holder voluntarily ceases market-making. In scenarios where the price crosses these preset boundaries, this constitutes a classic First Passage Problem, for which Darling and Siegert provided a derivation of the stopping time distribution using the Laplace Transform in 1953 [5]. We have further discovered that the expected value of the payoff under the risk-neutral measure contains exclusively terms of the form $\tau^{n}\exp(-r\tau)$ which are associated with the stopping time $\tau$. Therefore, we employ the Laplace Transform and its implications to derive the present value when the holder does not elect to terminate prematurely. In situations where the contract holder actively ceases market-making, we draw inspiration from the pricing methodology for Perpetual American Puts described by Shreve (2004) [6] and use numerical optimization to determine the optimal stopping strategy.
III Derivation
III-A Summary of Model Derivation
Uniswap V3 can be divided into two primary pricing components: the Liquidity Provider (LP) segment and the Rebate segment. The LP component is structured to accommodate either European or American option styles. Under the European construct, the Uniswap contract realizes its terminal value solely upon the price breaching the predetermined upper or lower thresholds. In contrast, the American framework allows the contract holder the flexibility to terminate the automated market-making process at any chosen point in time. The valuation of the Rebate segment hinges on the method selected by the contract holder for Rebate withdrawal. The upper limit scenario assumes the possibility of continuous Rebate extraction at the precise point of accrual. Meanwhile, the lower limit approach considers a singular, comprehensive withdrawal of all Rebates coinciding with the termination of the contract.
III-B Divide and Conquer
We initially assume that the ETHUSDT price follows a Geometric Brownian Motion (GBM) characterized by drift $\mu$ and annualized volatility $\sigma$ :
$dS(t)=\mu S(t)dt+\sigma S(t)dB_{t}$
or alternatively:
$S(t)=S_{0}\exp((\mu-\frac{\sigma^{2}}{2})t+\sigma B_{t})$
Define the unit price of ETHUSDT as follows
$P(t)=\frac{S(t)}{S_{0}}=\exp((\mu-\frac{\sigma^{2}}{2})t+\sigma B_{t})$
We partition the pricing structure of Uniswap V3 into two segments:
$V=V_{\text{LP}}+V_{\text{fee}}$
III-B1 Liquidity Provider segment
The unit price of the LP segment within the Uniswap V3 Contract, constrained by $S_{H}>S_{0}>S_{L}$, can be expressed as:
$V_{\text{LP}}(P_{t})$ =$=\begin{cases}L_{q}P(t)(\frac{1}{\sqrt{L}}-\frac{1}{\sqrt{H}}),P(t)<L\\ L_{q}(2\sqrt{P(t)}-\sqrt{L}-\frac{P(t)}{\sqrt{H}}),L<P(t)<H\\ L_{q}(\sqrt{H}-\sqrt{L}),H<P(t)\end{cases}$
where $H,L$ represent the unit bounds:
$H$ $=\frac{S_{H}}{S_{0}}$
$L$ $=\frac{S_{L}}{S_{0}}$
Additionally, the Liquidity Parameter is defined as:
$L_{q}=\frac{1}{2-\sqrt{L}-\frac{1}{\sqrt{H}}}$
III-B2 Rebate segment
The formulation for the rebate fee segment is comparatively straightforward:
$V_{\text{fee}}=C_{a}L_{q}V_{LP}(P_{t})t=365\times C_{d}L_{q}V_{LP}(P_{t})t$
where $C_{a},C_{d}$ are the daily and annual rebate rate.
III-C European LP Pricing
We are interested in evaluating:
$V(0)=\mathbb{E}(V(t))=\mathbb{E}(V_{\text{LP}}(t))+\mathbb{E}(V_{\text{fee}}(t))$
For the Uniswap V3 opsition, the stopping time $\tau$ is defined as:
$\tau=\inf\limits_{t<\infty}\{S(t)\leq S_{L}\text{ or }S(t)\geq S_{H}\}$
If the contract cannot be terminated before the stopping time, the current valuation of Uniswap V3 can be formulated as:
$V_{E}(0)$ $=\mathbb{E}(V(t))$
$=\mathbb{E}(V_{\text{LP}}(t))+\mathbb{E}(V_{\text{fee}}(t))$
$=\mathbb{E}[e^{-r\tau}V_{\text{LP}}(H)|S(\tau)=S_{H}]+\mathbb{E}[e^{-r\tau}V_{\text{LP}}(L)|S(\tau)=S_{L}]$
$+\mathbb{E}(V_{\text{fee}}(t))$
$=V_{\text{LP}}(H)\mathbb{E}[e^{-r\tau}|S(\tau)=S_{H}]+V_{\text{LP}}(L)\mathbb{E}[e^{-r\tau}|S(\tau)=S_{L}]$
$+\mathbb{E}(V_{\text{fee}}(t))$
The normalized log-unit price process of ETHUSDT is given by:
$\bar{W}(t)=\frac{\ln P(t)}{\sigma}=(\frac{\mu}{\sigma}-\frac{\sigma}{2})t+B_{t}$
Let us denote:
$\mu^{\prime}=\frac{\mu}{\sigma}-\frac{\sigma}{2}$
Thus, we have:
$\bar{W}(t)=x+\mu^{\prime}t+B_{t}$
where $x$ is the start point, $a,b$ are designated as the unit bounds:
$b$ $=\frac{\ln H}{\sigma},b^{\prime}=b-x$
$a$ $=\frac{\ln L}{\sigma},a^{\prime}=x-a$
Moreover, it also holds that:
$\tau=\inf\limits_{t<\infty}\{\bar{W}(t)\leq a\text{ or }\bar{W}(t)\geq b\}$
Utilizing the Laplace Transform, we can express the following expectations:
$\mathbb{E}[e^{-r\tau}|S(\tau)=S_{H}]=\exp({\mu^{\prime}b^{\prime}})\frac{\sinh[a^{\prime}\sqrt{\mu^{\prime 2}+2r}]}{\sinh[(b-a)\sqrt{\mu^{\prime 2}+2r}]}$
$\mathbb{E}[e^{-r\tau}|S(\tau)=S_{L}]=\exp({-\mu^{\prime}a^{\prime}})\frac{\sinh[(b^{\prime}\sqrt{\mu^{\prime 2}+2r}]}{\sinh[(b-a)\sqrt{\mu^{\prime 2}+2r}]}$
III-D American LP Pricing
In Euro-like pricing, liquidity providers can only exit when the price reaches the boundary. We further propose an American-like pricing model that allows liquidity providers to exit at any time. As Uniswap V3 contracts are perpetual, the termination of the contract relies exclusively on price movements, independent of time $t$. Therefore, it can be assumed that there exist boundaries determined solely by price levels. This implies that exit strategies should be based only on price considerations, analogous to the derivation of the perpetual American put option in Shreve’s work[6]. So we have:
$L<L_{1}<x<L_{2}<H$
where terminating automated market-making at prices $L_{1}$ or $L_{2}$. A set of stopping boundaries $L_{1},L_{2}$ can be determined through numerical optimization. First, we denote $c,d$ as previous section:
$d$ $=\frac{\ln L_{2}}{\sigma},d^{\prime}=d-x$
$c$ $=\frac{\ln L_{1}}{\sigma},c^{\prime}=x-c$
Then the equivalent price of the European contract:
$V_{E}$ $=LP_{L_{2}}\exp({\mu^{\prime}d^{\prime}})\frac{\sinh[c^{\prime}\sqrt{\mu^{\prime 2}+2r}]}{\sinh[(d-c)\sqrt{\mu^{\prime 2}+2r}]}$
$+LP_{L_{1}}\exp({-\mu^{\prime}c^{\prime}})\frac{\sinh[(d^{\prime}\sqrt{\mu^{\prime 2}+2r}]}{\sinh[(d-c)\sqrt{\mu^{\prime 2}+2r}]}$
The American contract can be priced using optimization techniques as following:
$V_{A}=\mathop{\max_{L<L_{1}<x<L_{2}<H}V_{E}(L_{1},L_{2})}$
III-E Rebate Pricing
For the rebate segment, if we assume that we can withdraw rebates continuously,(it is an upper bound of rebate)
[…]
$=C_{a}L_{q}V_{LP}(P_{t})(\cdot\mathbb{E}\left[\int_{0}^{\tau}e^{-rt}dt\right]$
$=C_{a}L_{q}V_{LP}(P_{t})\cdot\mathbb{E}\left[\frac{1-e^{-r\tau}}{r}\right]$
$=\frac{C_{a}L_{q}V_{LP}(P_{t})}{r}\cdot(1-\mathbb{E}\left[{e^{-r\tau}}\right])$
where,
$\mathbb{E}[e^{-r\tau}]:=F(r)$
$=\frac{e^{-\mu^{\prime}a^{\prime}}\operatorname{sh}\left(b^{\prime}\sqrt{2r+\mu^{\prime 2}}\right)+e^{\mu^{\prime}b^{\prime}}\operatorname{sh}\left(a^{\prime}\sqrt{2r+\mu^{\prime 2}}\right)}{\operatorname{sh}\left((b-a)\sqrt{2r+\mu^{\prime 2}}\right)}$
For the lower bound, we assume that we can withdraw all rebates when LPs close their position,
$\mathbb{E}(V_{\text{fee}}(t))=\mathbb{E}(C_{a}L_{q}V_{LP}(P_{t})e^{-r\tau}\tau)=C_{a}L_{q}V_{LP}(P_{t})\mathbb{E}(e^{-r\tau}\tau)$
where,
[…]
=$=\mathcal{L}\{\tau p(\tau)\}(r)$
$=-\frac{dF(r)}{dr}$
$=\frac{(b-a)(e^{-a^{\prime}\mu^{\prime}}\operatorname{sh}[b^{\prime}\sqrt{2r+\mu^{\prime 2}}]+e^{b^{\prime}\mu}\operatorname{sh}[a^{\prime}\sqrt{2r+\mu^{\prime 2}}])}{\operatorname{th}[(b-a)\sqrt{2r+\mu^{\prime 2}}]\operatorname{sh}[(b-a)\sqrt{2r+\mu^{\prime 2}}](\sqrt{2r+\mu^{\prime 2}})}$
$\ \ \ \ -\frac{b^{\prime}e^{-a^{\prime}\mu^{\prime}}\operatorname{ch}[b^{\prime}\sqrt{2r+\mu^{\prime 2}}]+ae^{b\mu}\operatorname{ch}[a^{\prime}\sqrt{2r+\mu^{\prime 2}}]}{\operatorname{sh}[(b-a)\sqrt{2r+\mu^{\prime 2}}](\sqrt{2r+\mu^{\prime 2}})}$
Figure 1: Relation between spot price and present value
Figure 2: Euro-Uniswap price distribution: Monte Carlo vs analytical solution
As illustrated in Figure 1, the blue line represents the variation of the payoff pricing model in response to changes in the spot price. The orange and green lines depict the corresponding numerical changes in the pricing model of a European option, where the orange line incorporates transaction costs and discounting, while the green line excludes these factors, both following the fluctuations in the spot price. We can observe that, under the current market parameters, if the option pricing model does not include transaction costs, it results in an expected loss at the current price, meaning that the present value is less than 1. Only by incorporating the potential transaction costs and their discounted value can we generate a positive return.
Figure 2 provides a Monte Carlo simulation to model the price dynamics of an asset following a Geometric Brownian Motion (GBM). The initial asset price $S_{0}$ was set to 1, with volatility $\sigma=0.6$, drift $\mu=0$, risk-free rate $r=0.04$, and boundary levels $H=1.2$ (upper bound) and $L=0.8$ (lower bound). We ran the simulation 10,000 times, calculating the payoff at each simulation based on whether the price reached the upper or lower boundary, then discounting the payoff to obtain the expected value. The Monte Carlo estimates closely matched the analytical solution derived from our new pricing formula, confirming the accuracy of our approach.
The final distribution of the outcomes, represented by a histogram, exhibited a bimodal shape. This reflects the nature of the problem, where the asset price is either stopped at the upper bound $H$ or the lower bound $L$ due to the stopping conditions inherent in the Uniswap V3 liquidity model. The two peaks correspond to the outcomes where the price reaches these boundaries, capturing the distinctive characteristics of liquidity provision in a decentralized exchange framework. This bimodal distribution highlights the risk/reward trade-offs faced by liquidity providers in such a system, where extreme price movements lead to clear-cut outcomes based on boundary interactions.
III-F Formulation of Case-Specific Models
Based on the derivations from the foundational section, we have obtained the pricing formulas for the following four scenarios. For the pricing formula of the European LP option problem under a GBM process, assuming that transaction fees can only be collected upon exiting the position, the pricing formula is as follows:
$V_{t}$
$=LP_{H}e^{\mu^{\prime}b^{\prime}}\frac{\operatorname{sh}(a^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((b-a)\sqrt{2r+\mu^{\prime 2}})}$
$+LP_{L}e^{\mu^{\prime}a^{\prime}}\frac{\operatorname{sh}(b^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((b-a)\sqrt{2r+\mu^{\prime 2}})}$
$+C_{a}L_{q}V_{LP}(P_{t})$
$(\frac{(b-a)(e^{\mu^{\prime}a^{\prime}}\operatorname{sh}(b^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}b^{\prime}}\operatorname{sh}(a^{\prime}\sqrt{2r+\mu^{\prime 2}}))}{\sqrt{2r+\mu^{\prime 2}}\operatorname{th}((b-a)\sqrt{2r+\mu^{\prime 2}})\operatorname{sh}((b-a)\sqrt{2r+\mu^{\prime 2}})}$
$-\frac{e^{\mu^{\prime}a^{\prime}}b^{\prime}\operatorname{ch}(b^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}b^{\prime}}a^{\prime}\operatorname{ch}(a^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\sqrt{2r+\mu^{\prime 2}}\operatorname{sh}((b-a)\sqrt{2r+\mu^{\prime 2}})})$
When we assume that transaction fees can be continuously collected during the lifetime of the position, the pricing formula becomes:
$V_{t}$ $=LP_{H}e^{\mu^{\prime}b^{\prime}}\frac{\operatorname{sh}(a^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((b-a)\sqrt{2r+\mu^{\prime 2}})}$
$+LP_{L}e^{\mu^{\prime}a^{\prime}}\frac{\operatorname{sh}(b^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((b-a)\sqrt{2r+\mu^{\prime 2}})}$
$+\frac{C_{a}L_{q}V_{LP}(P_{t})}{r}(1$
$-\frac{e^{\mu^{\prime}a^{\prime}}\operatorname{sh}(b^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}b^{\prime}}\operatorname{sh}(a^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((b-a)\sqrt{2r+\mu^{\prime 2}})})$
For the pricing formula of the American LP option problem, assuming that transaction fees can only be collected upon exiting the position, the pricing formula is as follows:
$V_{t}=LP_{L_{2}}e^{\mu^{\prime}d^{\prime}}\frac{\operatorname{sh}(c^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((d-c)\sqrt{2r+\mu^{\prime 2}})}$
$+LP_{L_{1}}e^{\mu^{\prime}c^{\prime}}\frac{\operatorname{sh}(bd^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((d-c)\sqrt{2r+\mu^{\prime 2}})}+C_{a}L_{q}V_{LP}(P_{t})($
$\frac{(d-c)(e^{\mu^{\prime}c^{\prime}}\operatorname{sh}(d^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}d^{\prime}}\operatorname{sh}(c^{\prime}\sqrt{2r+\mu^{\prime 2}}))}{\sqrt{2r+\mu^{\prime 2}}\operatorname{th}((d-c)\sqrt{2r+\mu^{\prime 2}})\operatorname{sh}((d-c)\sqrt{2r+\mu^{\prime 2}})}$
$-\frac{e^{\mu^{\prime}c^{\prime}}d^{\prime}\operatorname{ch}(d^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}d^{\prime}}c^{\prime}\operatorname{ch}(c^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\sqrt{2r+\mu^{\prime 2}}\operatorname{sh}((d-c)\sqrt{2r+\mu^{\prime 2}})})$
When we assume that transaction fees can be continuously collected during the lifetime of the position, the pricing formula for the American LP option problem becomes:
$V_{t}$ $=LP_{L_{2}}e^{\mu^{\prime}d^{\prime}}\frac{\operatorname{sh}(c^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((d-c)\sqrt{2r+\mu^{\prime 2}})}$
$+LP_{L_{1}}e^{\mu^{\prime}c^{\prime}}\frac{\operatorname{sh}(bd^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((d-c)\sqrt{2r+\mu^{\prime 2}})}$
$+\frac{C_{a}L_{q}V_{LP}(P_{t})}{r}(1$
$-\frac{e^{\mu^{\prime}c^{\prime}}\operatorname{sh}(d^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}d^{\prime}}\operatorname{sh}(c^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((d-c)\sqrt{2r+\mu^{\prime 2}})})$
IV Sensitivity Analysis
In the previous sections, we derived the analytical solution of Uniswap V3 positions. In this section, we focus on the sensitivity analysis of the Uniswap V3 position, specifically examining the Greeks associated with a typical European option framework. We previously derived the closed-form solutions for the pricing mechanisms of Uniswap V3 positions. Here, we extend our analysis to assess how variations in underlying parameters affect the pricing—commonly referred to as ”Greeks.”
TABLE I: $C=0.2$ $r=0.05$ $\sigma=0.7$
Model PV Delta Gamma Vega Rho
Payoff 1.000 0.561 -0.632 nan nan
European 1.422 0.427 -0.747 1.720 -25.892
American 1.696 0.539 -0.989 0.843 -19.425
TABLE II: $C=0.04$ $r=0.05$ $\sigma=0.4$
Model PV Delta Gamma Vega Rho
Payoff 1.000 0.784 -0.883 nan nan
European 0.643 0.074 0.073 1.296 -7.546
American 1.018 0.765 -1.059 0.200 -2.019
Figure 3: European LP Pricing Analysis
Our subsequent Greeks analysis will be conducted based on two sets of market parameters corresponding to different market conditions. One scenario features high transaction fees and high market volatility, referred to as ’good market situation’, as these conditions allow Uniswap V3 positions to accrue more fees. The other scenario involves low transaction fees and low market volatility, termed ’good market situation’.
Taking the European LP option pricing model as an example, we present the valuation discrepancies between pricing under good/bad market conditions and the payoff model at different price levels. The good/bad market conditions referenced here are based on the market parameters used in Table I and Table I. The range order price interval is set at (0.8,1.2). As shown in Figure 3, under the calculations based on the European LP option pricing model, there is a significant difference between the current Uniswap V3 payoff value and its actual present value. This is because the current value of a Uniswap V3 position is not limited to the extractable payoff; it also incorporates the discounted value of future fees that can still be collected if the position is not immediately withdrawn.
In scenarios with relatively low volatility and low transaction fees, the analysis shows that if the Uniswap V3 position is not withdrawn now, its future value will depreciate, and the discounted value of future fees will be insufficient to offset the loss caused by the position’s depreciation, as depicted by the green line in the Figure 3. Conversely, under good market conditions with higher volatility and transaction fees, the orange line in the Figure 3 demonstrates that the actual value of the Uniswap V3 position exceeds its payoff value.
Figure 4: European LP Delta Analysis
IV-A Delta Analysis
In historical Uniswap V3 trading practices, to mitigate the uncertainty of future returns, delta hedging is often employed to maintain delta-neutral risk exposure for Uniswap V3 positions. However, based on our European LP option pricing model, we find that delta risk within the $(L,H)$ range differs from the delta derived from payoff-based calculations.
Taking the comparison between the orange and blue lines in Figure 4 as an example (since it is uncommon to persist with Uniswap V3 positions under unfavorable market conditions), we observe that when the price is at 0.9, the delta value of the orange line exceeds that of the blue line. This indicates that the actual delta risk at this point is higher than the delta risk derived from payoff-based pricing. Consequently, more delta needs to be hedged to achieve delta neutrality. Conversely, when the price is at 1.1, the actual delta risk associated with holding the Uniswap V3 position is lower. This suggests that hedging excessive delta based on payoff-based calculations is unnecessary.
Figure 5: European LP Gamma Analysis
IV-B Gamma Analysis
By conducting a gamma analysis of the European LP pricing model versus the payoff-based pricing model, we can observe from Figure 5 that holding a Uniswap V3 position inherently involves gamma risk. Gamma risk refers to the situation where large gamma values cause significant fluctuations in delta with small changes in the underlying price $S$. In such scenarios, the delta hedging strategy becomes less effective, as even minor changes in $S$ can lead to substantial delta risk, necessitating more frequent adjustments to maintain delta neutrality. Conversely, when gamma values are small, delta remains relatively stable, reducing the need for frequent hedging.
As shown in the Figure 5, there is a significant difference between the gamma risk values derived from payoff-based modeling and those from the European LP pricing model. This highlights the higher requirements for managing gamma risk in Uniswap V3 positions, emphasizing the need for more refined risk control strategies.
IV-C Vega Analysis
Vega measures the sensitivity of the valuation of a Uniswap V3 European pricing position to changes in volatility. From the Figure 6, we observe the variation of vega with respect to spot prices under different market conditions. Notably, there is no vega estimation graph for the payoff-based pricing model, as this model assumes no correlation between price and volatility, and thus does not account for vega risk.
From the Figure 6, it is evident that vega remains consistently negative within the defined $(L,H)$ range. This indicates that holding a Uniswap V3 position inherently represents a short volatility strategy, implying that lower volatility leads to higher returns.
Figure 6: European LP Vega Analysis
IV-D Greeks Comparison between different range order
After analyzing the unique patterns exhibited by the Greeks under Uniswap V3 positions, we aim to compare the Greeks of Uniswap V3 positions established with different $(L,H)$ ranges to uncover the relationship between Uniswap V3 pricing and the choice of $(L,H)$. The results are illustrated in the figure.
We set up three groups of Uniswap V3 range orders with price ranges of $(0.8,1.1)$, $(0.8,1.2)$, and $(0.8,1.3)$. The assumed market conditions are $r=0.05$, $C=0.2$, $\mu=0$, and $\sigma=0.7$. From Figure 7 ,we observe that as the pricing range widens, the European LP pricing model’s valuation results tend to increase. This is because a wider pricing range implies that the asset is expected to remain within the pricing range for a longer period, thereby extending the expected time for fee collection.
Figure 7: Different Range LP Pricing Comparison
An interesting observation arises from the delta comparison in Figure 8 . We find that as the pricing range widens, the delta values generally increase. This is because when the pricing range is very wide, the proportional impact of price fluctuations on the exchange between the underlying asset and currency diminishes relative to the total position value. Consequently, the position value remains strongly correlated with the underlying asset price, resulting in a higher delta. Conversely, when the pricing range is narrow, even small price movements can trigger significant exchange activity, weakening the correlation between position value and underlying asset price, thus reducing delta.
Figure 8: Different Range LP Pricing delta Comparison
Although wider pricing ranges lead to larger delta risks, the absolute value of gamma decreases, as shown in Figure 9. This indicates that delta hedging becomes easier for wider pricing ranges compared to narrower ones. With a broader pricing range, delta changes more smoothly as the spot price fluctuates, simplifying the delta hedging process.
Figure 9: Different Range LP Pricing gamma Comparison
Finally, as demonstrated in Figure 10, even with changes to the pricing range, vega remains strictly negative. This confirms that the relationship between price and volatility consistently exhibits a negative correlation.
Figure 10: Different Range LP Pricing Pricing Analyiss versus vol
To validate the effectiveness of our Greeks, we need to perform numerical tests. Specifically, based on the Taylor expansion method, the following equation holds:
$\Delta V\approx\text{Delta}\times\Delta S+\frac{1}{2}\text{Gamma}\times\Delta S^{2}+\text{Vega}\times\Delta\sigma$
We can use two sets of different parameters, $S$ and $\sigma$, to calculate the valuation change of our Uniswap V3 position before and after market fluctuations (based on the present value derived from our pricing model). If the Greeks we calculate, when integrated into the right-hand side of the equation, can estimate the pricing with an acceptable level of accuracy, it will sufficiently demonstrate the validity of our Greeks.
The parameters we set are as follows, for the market before the fluctuation:
$C=0.2,r=0.05,\mu=0,\sigma=0.7$
$S=1,H=1.1,L=0.8$
Taking ETH as an example, under general circumstances, we assume a 5 % decrease in price and a 10 % increase in volatility itself. Using our pricing model, the valuation of the Uniswap V3 position before and after the market fluctuation is calculated as follows:
The Current Pv is: 1.02110809997664861.0211080999766486 1.02110809997664861.0211080999766486
The new PV is: 0.99304363754849550.9930436375484955 0.99304363754849550.9930436375484955
The Position Changing is $-0.028064462428153192$
The Greeks obtained through calculation, along with the corresponding Taylor expansion results, are as follows:
The Current delta is: 0.197906601375752440.19790660137575244 0.197906601375752440.19790660137575244
The Current gamma is: $-11.101273578173476$
The Current vega is: $-0.15958993211606654$
Taylor Expansion estimation is: $-0.034943217289629125$
The Current residential is: $-0.006878754861475933$
Thus, under this level of market changes, the pricing estimation derived from the Taylor expansion exhibits an approximate error of 0.6% compared to the actual valuation of the Uniswap V3 position.
IV-E The Impact of Volatility
The relationship between the volatility of the underlying asset and the pricing dynamics of the Uniswap V3 position is integral to our sensitivity analysis. Intuitively, we understand that according to the European option valuation model, the present value (PV) pricing should exhibit a monotonically negative correlation with volatility. This stems from the Automatic Market Maker (AMM) mechanism of Uniswap V3, which fundamentally behaves as a short volatility strategy. As the volatility increases, we anticipate a decrease in the valuation, signifying an inverse correlation.
However, the observed orange line in Figure 11deviates from this predicted relationship due to the dual pricing model for transaction fees, which has separate methodologies for calculating both upper and lower bounds. The rationale behind the lower bound is driven by the expectation of time until a transaction fee is collected, conceptualized through the theory of stopping times, which is directly proportional to volatility. Conversely, the upper bound assumes continuous trading throughout the transactional process, operating under the premise that serial transactions yield continuous fees, justifying the positive valuation associated with low volatility events.
We selected points with relatively low volatility, such as $\sigma$=0.02, to compare the pricing models of the two transaction fee calculation methods in terms of their present value (PV), as shown in Figure 12. In the PV-spot graph, the pricing model based on the upper bound of transaction fees consistently yields a higher value than the pricing model based on the lower bound of transaction fees.
Figure 11: relation between pv and $\sigma$ using different rebate calculation
Figure 12: relationship of pv-spot with different fee calculation
V future work
V-1 Hedging Strategies
Our pricing model is equipped to offer more accurate calculations of Greek risk measures such as delta, gamma, and vega. These calculations are crucial for developing sophisticated hedging strategies. By leveraging these measures, we can devise fully hedged strategies that maximize profitability while minimizing risk. Such strategies will not only benefit LPs by enhancing their returns and reducing risk but also contribute to the overall stability and efficiency of the DeFi ecosystem.
V-2 Adaptation to Uniswap v4
With the introduction of Uniswap v4, various new decentralized exchange (DEX) paradigms will be implemented, such as dynamic fee structures [7], which make $C$ a function of volatility. This simplification of the model allows LPs to adjust their positions less frequently in response to market conditions, making participation more accessible and efficient. Beyond dynamic fees, many potential enhancements and new models may emerge with Uniswap v4, requiring further adaptations and improvements to our frameworks.
V-3 Implied volatility
Our model is not only applicable for pricing but also serves as a tool for estimating implied volatility. Similar to the Black-Scholes-Merton (BSM) framework, we can infer the views of LPs on volatility through their positions. By aggregating these inferred volatilities, we can construct a liquidity surface that reflects the weighted perspectives of LPs.